Purdue University MA453 Intro to Algebra Spring 2011 Syllabus Instructor: Uli Walther Office: 746 MATH Phone: 49-41959

                                E-mail: walther@math.purdue.edu

1 General Information

1.1 Class time / location:

TTh 9:00-10:15pm, REC121.

1.2 Office Hours:

Tue 2:30-3:00 and Th 1:30-2:30.

1.3 Textbook:

Contemporary Abstract Algebra Joseph Gallian 6th edition

1.4 Webpage:

Under http://www.math.purdue.edu/~walther/teach you find this syllabus

as well as some other info concerning the course, including a link to our Project Rhea (see below).

1.5 Statement for Students with Disabilities

See the webpage, http://www.math.purdue.edu/~walther/teach.

1.6 Calculators

You may use NO calculators on the tests. Furthermore, in all cases, numerical answers alone will get you no credit. This class emphasizes understanding the process that leads to a solution, not the solution itself. Credit will be given for the explanation of the method (if correct) and how it is carried out.

2 Course Structure

If you disagree with the grading in any category, you need to resubmit the item in question with a written explanation stating why (and not just that) you deserve more credit.

2.1 Homework:

Homework will be collected every Thursday in class or in my office no later than 3pm. No late assignments will be accepted. Neither will homework deposited anywhere else.

The homework due in any given week is the homework corresponding to the material of the previous week. Homework must be readable and must be stapled. Illegible scribblings will receive no credit from the grader.

You are encouraged to attempt all the questions and discuss with your classmates. However, the write-up must be of your own.

• WORK MUST BE SHOWN TO EXPLAIN YOUR ANSWERS. NO CREDIT WILL BE GIVEN FOR JUST THE NUMERICAL AN-SWERS. THIS ALSO APPLIES TO EXAMS.

2.2 Project Rhea

You may earn up to 5% extra credit through writing essays on Rhea. There are suggested topics posted, but I am open to your suggestions.

2.3 Midterms:

There will be one midterm, Thursday February 24. Makeup tests may be given in extraordinary instances, but only with documented reasons.

2.4 Final Exam:

The Final Exam will be given during the Final Exam Week.

The location and date of the final will be announced in class.

2.5 Course Grade:

Your course grade will be determined using the following distributions:

Rhea  up to 5%
HW  30%
Midterm 30%
Final  40%